OpenStudy (anonymous):

if x^2+3x-70=(x+9)(x+b) , then a+b equals to? answers a) -17 B)-3 C)3 D)17 E)32

OpenStudy (anonymous):

a=7 and b=-10 then a+b=-3 B)

OpenStudy (anonymous):

B

myininaya (myininaya):

how did you guys get a where is a and how dose 9*(-10)=-70? this doesn't make any sense

myininaya (myininaya):

does*

myininaya (myininaya):

it might make alittle more since if she had said if x^2+3x-70=(x+a)(x+b), then a+b equals to? a=-7 and b=10 a+b=-7+10=3

myininaya (myininaya):

the answer is C

myininaya (myininaya):

judith do you understand?

myininaya (myininaya):

i guess not lol

OpenStudy (anonymous):

lol yeaa i get it thanks(: